# Modified Axelrod Model Showing Opinion Convergence And Polarization In Realistic Scale-Free Networks

### v1

@@ V Reducing Opinion Polarization By Introducing Sympathetic Agents
Here, we investigate the opinion evolution of systems with sympathetic agents in two ways. Figure 7 depicts the opinion distribution evolution when sympathetic agents are assigned in the beginning. As expected, the higher the value of $p_{\text{s}}$, the smaller the deviation in the attitudes of all agents. Although a sizable number of extremists is still present, a significant fraction of moderates remain. Nevertheless, this effect is significant only when $p_{\text{s}}\gtrsim 15\%$. Such a high percentage of sympathetic agents is very hard if not impossible to achieve in reality.
@@
A more pressing issue is how to reduce society polarization in an already polarized society. Therefore, we study what happens if we change a fixed portion of agents from “normal” to sympathetic after the society has evolved almost to its final state. More precisely, we first let the system evolve with $t=200$ using parameters listed in Table 1 so that it has almost equilibrated. Then, we will randomly pick and fix $p_{\text{s}}$ of the agents and make them sympathetic. Finally, we further evolve the system for an additional $t$ up to $400$. From Figure 8, we observe that one can significantly reduce opinion polarization only when $p_{\text{s}}\gtrsim 60\%$, a requirement that is even harder to achieve.
@@ VI Conclusions
In this paper, we have identified deficiencies in the Axelrod model of opinion dynamics and proposed a modified one as an attempt to build a more sophisticated model to describe opinion dynamics in human society. Major modifications include using scale-free networks replacing the static two-dimensional lattice grid together with a new set of interaction rules for similar and dissimilar agents. Our simulation results show that over essentially all reasonable parameters agents’ opinions on all features eventually become more and more polarized. Finally, we tried to reduce this extreme opinion polarization by introducing sympathetic agents that are open to talking and actively interacting with everyone. Their presence reduces opinion polarization but only when they appear en masse.

### v2

@@ V.1 The Case Of Unbiased Empathetic Agent Selection
We first consider the case when $\rho_{\text{s}}=1$ so that each agent has equally likely chance of chosen to be empathetic. Here, we investigate the opinion evolution of systems with these empathetic agents in two ways. Fig. 8 depicts the opinion distribution evolution when empathetic agents are assigned in the beginning. As expected, the higher the value of $p_{\text{s}}$, the smaller the deviation in the attitudes of all agents. Although a sizable number of extremists is still present, a significant fraction of moderates remain. Nevertheless, this effect is significant only when $p_{\text{s}}\gtrsim 35\%$. Such a high percentage of empathetic agents is very hard if not impossible to reach in reality.
@@
A more pressing issue is how to reduce society polarization in an already polarized society. Therefore, we study what happens if we change a fixed portion of agents from “normal” to empathetic after the society has evolved almost to its final highly polarized state. More precisely, we first let the system evolve with $t=200$ using parameters listed in Table 1 so that it has almost equilibrated. Then, we will randomly pick and fix $p_{\text{s}}$ of the agents and make them empathetic. Finally, we further evolve the system for an additional $t$ up to $400$. From Fig. 9, we observe that one can significantly reduce opinion polarization only when $p_{\text{s}}\gtrsim 40\%$, a goal that is even harder to attain.
V.2 The Case When Empathetic Agents Are Picked From The High Degree Nodes
In view of the ineffectiveness of picking empathetic agents randomly from the population, we consider the case of choosing them only from highly connected agents. In the real world, this corresponds to the case that a sizable number of social media influencers are empathetic. First we redo our simulations for the case of assigning empathetic agents right from the start by fixing $\rho_{\text{s}}=0.3$ as an illustration. As shown in Fig. 10, a significant fraction of moderates remains when $p_{\text{s}}\gtrsim 20\%$ (which is much less than the critical value of $p_{s}\approx 35\%$ for the case when $\rho_{\text{s}}=1$). The much more interesting case is when selected “normal” agents turn empathetic after the system has evolved with $t=200$. Fig. 11 depicts that system becomes much less polarized if $\rho_{\text{s}}=50\%$ and $p_{\text{s}}\gtrsim 30\%$. This corresponds to changing the behavior of about $60\%$ of those agents whose connectivity is above average. Furthermore, if picking $\rho_{\text{s}}=40\%$ and $p_{\text{s}}=60\%$, moderate opinion agents domain the system in the end. Compared to the case of randomly assigning empathetic agents, these strategies are much more viable although it is still highly challenging to implement.
@@ VI Conclusions
In this paper, we have identified deficiencies in the Axelrod model of opinion dynamics and proposed a modified one as an attempt to build a more sophisticated model to describe opinion dynamics in human society. Major modifications include using scale-free networks replacing the static two-dimensional lattice grid together with a new set of interaction rules for similar and dissimilar agents. Our simulation results show that over essentially all reasonable parameters agents’ opinions on all features eventually become more and more polarized. Besides, one or two dominating opinion clusters are formed together with numerous number of small opinion clusters. This is not unlike the big opinion differences between those living in cities and in rural areas as well as the emergence of political extremism due to distrust of the elites in the real world. We also discovered scaling in the distribution of the relative opinion cluster size $n(s,N)/N$. For B-A network, it follows a power law when $s\lesssim 10$ and switches to exponential decay when $15\lesssim s\lesssim 35$. Finally, we tried to reduce this extreme opinion polarization by introducing empathetic agents that are open to talking and actively interacting with their neighbors. Their presence reduces opinion polarization but only when they appear en masse. Our study suggests that one effective and relatively economical but highly challenging way to lessen the difference in an already highly polarized opinion population in the real world is to change the behavior of a significant portion of social media influencers making them more empathetic.